OpenStudy (anonymous):

An initial deposit of $1000 earns 7% interest, compounded monthly. How much will be in the account in 2½ years?

OpenStudy (anonymous):

amount = 1000( 1 + .07/12)^(12*2.5)

OpenStudy (anonymous):

okay, but I need to know the exact value to the nearest cent. I got $2680.73. But the thing is, it was wrong

OpenStudy (anonymous):

= 1000(1 + 0.0058333)^30 = 1190.64 dollars

OpenStudy (anonymous):

the formula is P = Po ( 1 + r/n)^nt

OpenStudy (anonymous):

i'll chec k it out on wolfram in case i made a mistake

OpenStudy (anonymous):

yup - thats correct

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

COOL

OpenStudy (anonymous):

Why did it count it wrong then?